media%2Fa77%2Fa771601d-1af2-4cde-99a9-ce
Answer #1

The object distance u, image distance v and focal length of a lens are related by the expression

\frac{1}{f} = \frac{1}{v}+ \frac{1}{u}

for the objective, \frac{1}{f_1} = \frac{1}{v_1}+ \frac{1}{u_1}

=> \frac{1}{20} = \frac{1}{v_1}+ \frac{1}{30}

=> v1 = 60 cm

and magnification m1 = -v1/u1 = - 60/20 = -3

distance between the objective and eyepiece is 66 cm

therefore, u2 = 66 - 60 = 6 cm

so, \frac{1}{20} = \frac{1}{v_2}+ \frac{1}{6}

=> v2 = - 8.57 cm

and so the magnification will be: m2 = -(-8.57)/6 = 1.43

and therefore the total magnification of the system is: m = m1m2 = -3(1.43) = - 4.29.

the final image will be virtual, magnified and upright.

The near point of a human eye is about 25 cm which basically means that within 25 cm from the eye, we can bring the image into focus and see a sharp image. Since this image due to the eyepiece is just 8.57 cm from the eye, therefore the scientist will suffer discomfort if viewing for prolonged period of time.